OpenStudy (rainbow_rocks03):

Evaluate. 5 • [3 + (16 – 4)] ÷ 3 A. 9 B. 19 C. 21 D. 25

OpenStudy (rainbow_rocks03):

@CallMeKiki

OpenStudy (anonymous):

D 25

OpenStudy (rainbow_rocks03):

no direct answer I want the explanation too

OpenStudy (anonymous):

ok so use PEMDAS Do the work in the () first 16-4=12

OpenStudy (anonymous):

then do 12+3=15

OpenStudy (rainbow_rocks03):

oh

OpenStudy (anonymous):

then times that by 5 15*5=75 then divide that by 3 75/3=25

OpenStudy (puppylover0617):

first do 16-4=12 then 12+3=15 then 15*5=75 last 75/3= 25

OpenStudy (rainbow_rocks03):

oh thx @nerd_3.14

OpenStudy (anonymous):

np

OpenStudy (callmekiki):

C: too fast ahaha